OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) -1 -5 0 -1 1 3 g(x) g(x) = 4x + 2 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer.

OpenStudy (anonymous):

@rishavraj @gorv @geekfromthefutur

OpenStudy (anonymous):

@awesomeness812 @KamiBug @leonardo0430

OpenStudy (gorv):

can u type x and f(x) again

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In a certain town, the inches of rainfall, x, on a certain day is described by two statements: If 3 times the rainfall is increased by 2, the rainfall is still less than 26 inches. 5 plus twice the rainfall is greater than 11 inches. Part A: Create a compound inequality to represent the rainfall range. (3 points) Part B: Can the rainfall in this town be 4 inches? Justify your answer by solving the inequalities in Part C: The average rainfall in another town is 5 inches, but the actual rainfall is within 4 inches of the average. Write and solve an inequality to find the range of rainfall in the town

OpenStudy (anonymous):

x f(x) -1 -5 0 -1 1 3

OpenStudy (anonymous):

can u guys please help

OpenStudy (gorv):

ok first we need to find f(x)

OpenStudy (anonymous):

ok

OpenStudy (gorv):

let f(x)=y y=slope*x+c this is the standard equation

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

f(x) goes by +4 each time as in -5 +4= -1 + 4 = 3

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so now i need help with part a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

??

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can u help me ?????

OpenStudy (anonymous):

me yea

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

y = 2x + 30 slope (32 -30)/(1-0) = 2 when x = 0, y = 30

OpenStudy (anonymous):

ok... for any linear equation the slope intercept form is y = mx + b m is the slope of the line b is the y-intercept, or where the line cuts the y axis so f(x) has a slope of 2 g(x) has a slope of 8 the larger the number for the slope, the steeper the line. -25 30 f(x) g(x) Reply Using Drawing so the questions... so to answer question 1. you can say, the point slope form of a line is y = mx + b. f(x) has a slope of 2 and g(x) has a slope of 8, g(x) is the line with the steepest slope.

OpenStudy (gorv):

we need to find slope

OpenStudy (anonymous):

f(x) has a y intercept of 30 and g(x) has a y intercept of -25... so f(x) has the greater intercept. hope this helps

OpenStudy (gorv):

slope =(f(x1)-f(x2))/(x2-x1)

OpenStudy (gorv):

f(x1)=-5 f(x2)=-1

OpenStudy (anonymous):

:)

OpenStudy (gorv):

x1=-1 x2=0

OpenStudy (anonymous):

so which one has the greater intercept ????/

OpenStudy (anonymous):

nvmmm i see now thanks :))))

OpenStudy (anonymous):

i have another problem @awesomeness812

OpenStudy (gorv):

g(x) has greater intercept

OpenStudy (anonymous):

yw\ea

OpenStudy (anonymous):

k